OpenStudy (anonymous):

solve. (5-y) + (6-y) - (5-y)=0

OpenStudy (anonymous):

5 - y + 6 - y - 5 + y = ???

OpenStudy (anonymous):

Or: \((5-5 + 6) + (y - y - y)\) = ???

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

and y - y -y = ???

OpenStudy (anonymous):

-y?

OpenStudy (anonymous):

Yes.. So 6 - y

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

6 - y = 0 add y both the sides..

OpenStudy (anonymous):

ok got it

OpenStudy (anonymous):

Welcome dear..